OpenStudy (anonymous):

for nonzero numbers w, x, and y, if x is twice y and y is (1/3) of w, what is teh ratio of w^2 to x^2

OpenStudy (anonymous):

2 / (1/3) = 6 6^2 = 36 therefore the ratio is 1:36

OpenStudy (anonymous):

x = 2y and y = (1/3)w 3y = w we have: w^2/x^2 = (3y)^2 / (2y)^2 = 9/4 - the y^2 cancels

OpenStudy (anonymous):

thanks